OpenStudy (anonymous):

Factoring Trinomials by Grouping. Does anyone know how to solve one of these problems? k^2 + 2k + kt + 2t Thank you!

OpenStudy (anonymous):

=k(k+2)+t(k+2) =(k+t)(k+2)

OpenStudy (anonymous):

So you would just group the two terms then the other two then factor out? I tried to factor out k from the first group then t from the second but after regrouping the outside terms, the two groups were not the same? I am probably now explaining this correctly.